Hi I think this is the right subject, Its called "Equations with like terms" And i have to combine them into one expression
the first one is
3x+4x+3= -39

Answer by TimothyLamb(4379)

---
3x + 4x + 3 = -39
3x + 4x + 3 - 3 = -39 - 3
3x + 4x = -42
7x = -42
7x/7 = -42/7
x = -6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php